# Radial amplitude equations for fully localised planar patterns

### v1

@@ 3.2 Other fully localised patterns
We now consider the other dihedral patterns $u_{H}$, $u_{R}$ and $u_{Q}$ defined in § 2.2. We note that the quadratic projections in (2.7), (2.9) & (2.11) all include $J_{n}(r)$ terms; this means that the quadratic terms in (3) partially lie in the kernel of $(1+\Delta_{n})^{2}$ and thus cannot be removed by normal form transformations.
To circumvent this problem, we take $\nu:=\varepsilon\hat{\nu}$ such that (3) becomes
@@
and $v_{n}^{(0)}=v_{n}^{(1)}=0$ for all $n\in\mathbb{Z}$ with $3\nmid n$, such that (3.6c) becomes
@@
for some function $f$ of irrational frequencies $\{\sqrt{3}\,r,\sqrt{7}\,r\}$. We restrict to the primary frequency $J_{3n}(r)$ and define $A_{h}:=\frac{1}{3}A$, so that we obtain the amplitude equation
@@
and $v_{n}^{(0)}=v_{n}^{(1)}=0$ for all $n\in\mathbb{Z}$ with $3\nmid n$, where $A_{q}$ satisfies
@@
Stationary, real solutions of equations (3.8) and (3.11) can be investigated using phase plane analysis. The standard set up is that there is a small amplitude homoclinic solution that bifurcates off the trivial state and broadens to a “table-top” localised pattern as one approaches the Maxwell point, $\hat{\mu}=\hat{\mu}_{M}$; see Figure 4.
For general equations of the form
@@ 4 Fully localised hexagons in reaction–diffusion systems
so that (4.3) becomes
@@
The $\mathcal{O}(\varepsilon^{2})$ terms are hence absorbed into the $\mathcal{O}(\varepsilon^{4})$ equation, which does not affect our analysis.
@@
where we have defined $\mathbf{Q}_{0}:=\mathbf{Q}(\hat{U}_{0},\hat{U}_{0})$ and $\mathbf{C}_{0}:=\mathbf{C}(\hat{U}_{0},\hat{U}_{0},\hat{U}_{0})$. Again the $\mathcal{O}(\varepsilon)$ terms are absorbed into the $\mathcal{O}(\varepsilon^{4})$ equation. We take the inner product of $(\mathbf{M}_{1}-\Delta_{3n})\mathbf{u}_{3n}^{(2)}$ with $\hat{U}_{1}^{*}$ and restrict to the prime frequency $J_{3n}(k_{c}r)$, resulting in
@@ 5 Conclusion
Our results raise new questions regarding the connections between fully localised patterns and the localised axisymmetric patterns studied in [15, 18]. In the Galerkin finite-Fourier approach of Hill et al. [8, 9], it was found that some localised dihedral patterns (including hexagons) are qualitatively similar to axisymmetric spots, while others were instead related to axisymmetric rings. We note, however, that the approach described in this paper does not work for purely axisymmetric patterns, since we do not have any equivalent identities for dealing with nonlinear products of just $J_{0}$ Bessel functions. Furthermore, it is not clear that one would be able to recover our results in a Galerkin finite-Fourier approximation, since we again lose the nonlinear identities that are vital to our analysis.
@@ Acknowledgements
DJH gratefully acknowledges support from the Alexander von Humboldt Foundation. The idea for this work began during a Postdoc/PhD seminar at Institut Mittag–Leffler during the research program ‘Order and Randomness in Partial Differential Equations’, supported by the Swedish Research Council under grant no. 2021-06594 in the winter semester of 2023.

### v2

@@ 3.2 Other fully localised patterns
We now consider the other dihedral patterns $u_{H}$, $u_{R}$ and $u_{Q}$ defined in § 2.2. We note that the quadratic projections in (2.11), (2.13) & (2.15) all include $J_{n}(r)$ terms; this means that the quadratic terms in (3.1) do not lie entirely in the range of $(1+\Delta_{n})^{2}$ and thus cannot be removed by normal form transformations. We note that this is not unique to our radial projections, and is also observed for hexagons in Cartesian coordinates due to the presence of resonant triads in Fourier space (see Section 5.4 in [10], for example).
To circumvent this problem, we take $\nu:=\varepsilon\hat{\nu}$ such that (3.1) becomes
@@
and $v_{n}^{(0)}=v_{n}^{(1)}=0$ for all $n\in\mathbb{Z}$ with $n$ not a multiple of $3$, such that (3.6c) becomes
@@
for some function $f$ of irrational frequencies $\{\sqrt{3}\,r,\sqrt{7}\,r\}$. We define $A_{h}:=\frac{1}{3}A$ and again obtain a solvability condition for $A_{h}$ in the form of the amplitude equation
@@
and $v_{n}^{(0)}=v_{n}^{(1)}=0$ for all $n\in\mathbb{Z}$ with $n$ not a multiple of $3$, where $A_{q}$ satisfies
@@
Since (3.11) is independent of $b_{n}$, it follows from Remark 2.8 that (3.11) is also the amplitude equation for the superposition of two arbitrarily rotated hexagon lattices $u_{\alpha}$ for any rotation angle $\alpha\in(0,\frac{\pi}{6})$. In order to study steady localised solutions of (3.8) and (3.11), we again restrict to the real invariant subspaces $A_{h}\in\mathbb{R}$, $A_{q}\in\mathbb{R}$. Stationary, real solutions of equations (3.8) and (3.11) can be investigated using phase plane analysis, where we observe a small amplitude homoclinic solution that bifurcates off the trivial state and broadens to a “table-top” localised pattern as one approaches the Maxwell point, $\hat{\mu}=\hat{\mu}_{M}$; see Figure 4. In particular, for general equations of the form
@@ 4 Fully localised hexagons in reaction–diffusion systems
so that $\Delta_{3n}\mathbf{u}_{3n}$ becomes $\Delta_{3n}\mathbf{u}_{3n}+\varepsilon(\mathcal{D}_{3n}+\mathcal{D}_{-3n})\hat{\mathcal{D}}_{0}\mathbf{u}_{3n}+\varepsilon^{2}\hat{\Delta}_{0}\mathbf{u}_{3n}$ and (4.3) becomes
@@
Here we have used the fact that $\Delta_{n}=\Delta_{m}+\frac{(m^{2}-n^{2})}{r^{2}}$ and $(\Delta_{3n\mp 1}+k_{c}^{2})\mathcal{D}_{\pm 3n}J_{3n}(k_{c}r)=0$, and so
$\begin{split}(\Delta_{3n}+k_{c}^{2})\left(\mathcal{D}_{3n}+\mathcal{D}_{-3n}\right)J_{3n}(k_{c}\,r)={}&k_{c}(\Delta_{3n}+k_{c}^{2})J_{3n-1}(k_{c}\,r)-k_{c}(\Delta_{3n}+k_{c}^{2})J_{3n+1}(k_{c}\,r)\\ ={}&\frac{1}{r^{2}}\left[\left(\mathcal{D}_{3n}+\mathcal{D}_{-3n}\right)J_{3n}(k_{c}r)-\frac{(6n)^{2}}{r}J_{3n}(k_{c}r)\right]\end{split}$
Evaluating on the slow spatial scale $R$ we conclude that $(\Delta_{3n}+k_{c}^{2})\left(\mathcal{D}_{3n}+\mathcal{D}_{-3n}\right)J_{3n}(k_{c}\,r)=\mathcal{O}(\varepsilon^{2})$, since $\frac{1}{r^{2}}=\varepsilon^{2}\,\frac{1}{R^{2}}$, and hence this term is absorbed into the $\mathcal{O}(\varepsilon^{4})$ equation, which does not affect our analysis.
@@
where we have defined $\mathbf{Q}_{0}:=\mathbf{Q}(\hat{U}_{0},\hat{U}_{0})$ and $\mathbf{C}_{0}:=\mathbf{C}(\hat{U}_{0},\hat{U}_{0},\hat{U}_{0})$. Again the $\mathcal{O}(\varepsilon)$ terms are absorbed into the $\mathcal{O}(\varepsilon^{4})$ equation. We take the inner product of $(\mathbf{M}_{1}-\Delta_{3n})\mathbf{u}_{3n}^{(2)}$ with $\hat{U}_{1}^{*}$ and obtain the following solvability condition
@@ 5 Conclusion
Our results raise new questions regarding the connections between fully localised patterns and the localised axisymmetric patterns studied in [15, 18]. In the Galerkin finite-Fourier approach of Hill et al. [8, 9], it was found that some localised dihedral patterns (including hexagons) are qualitatively similar to axisymmetric spots, while others were instead related to axisymmetric rings. We note, however, that the approach described in this paper does not work for purely axisymmetric patterns, since we do not have any equivalent identities for dealing with nonlinear products of just $J_{0}$ Bessel functions. Furthermore, it is not clear that one would be able to recover our results in a Galerkin finite-Fourier approximation, since we again lose the nonlinear identities that are vital to our analysis. Notably, Remark 2.8 tells us that arbitrarily rotated hexagons possess the amplitude equation (3.11) for all rotations $\alpha\in(0.\frac{\pi}{6})$, suggesting that the quadratic-cubic Ginzburg–Landau equation seen in (3.8) and (3.11) is highly generic.
@@ Acknowledgements
DJH gratefully acknowledges support from the Alexander von Humboldt Foundation. The idea for this work began during a Postdoc/PhD seminar by Bastian Hilder and Jonas Jansen at Institut Mittag–Leffler during the research program ‘Order and Randomness in Partial Differential Equations’, supported by the Swedish Research Council under grant no. 2021-06594 in the winter semester of 2023.